OpenStudy (anonymous):

Solve x + 7 = -2(4x + 10).

OpenStudy (hoblos):

x + 7 = -2(4x + 10) x+7 = -8x - 20 add(8x) and subtract (7) 9x = -27 divide by 9 x = -3